OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14262 — var-5-6k
Incident Report for Network & Infrastructure
Resolved
There's a hardware issue with a TCAM on this component.
We're working to fix the issue.

Update(s):

Date: 2015-08-10 09:33:53 UTC
TPIX has performed some actions on their side. We've re-enabled and are closely monitoring the traffic.

Date: 2015-08-05 10:35:58 UTC
The issue originated from TPIX.

[snip]
We still have some problems with switch. We will do reboot within 30 minutes. After that we will inform you about status.
We are very sorry about it.
[\\snip]

Waiting for confirmation from TPIX before remounting the traffic.

Date: 2015-08-05 10:34:37 UTC
We corrected the hardware issue. The TPIX peer IPs are still not working. We've cut off sessions with the TPIX peers to regain connectivity.
Posted Aug 05, 2015 - 08:24 UTC